# Uncovering the first-infall history of the LMC through its dynamical impact in the Milky Way halo

### v1

@@ 6 Conclusions
  • We compare the mean latitudinal velocities in the heliocentric frame of $\sim 1,200$ RR Lyrae stars located within a Galactocentric radius of 50-100 kpc with those from the mock star catalogs generated by our 50 MW-LMC simulations. We observe that the RR Lyrae sample exhibits a mean latitudinal velocity of $\langle v_{b}\rangle=30.8\pm 4.0$ km/s.
  • This result aligns with the outcomes ($\langle v_{b}\rangle=28\pm 5$ km/s) from simulations in which the LMC follows a first-infall orbital scenario and the MW halo is spherical ($q=1.0$). In contrast, the results from simulations where the LMC has a second-passage orbital scenario are consistently lower than $19\pm 5$ km/s for various parameter combinations. Therefore, our comparative study supports the first-infall scenario for the LMC.

### v2

@@ 6 Conclusions
  • We compare the mean latitudinal velocities in the Galactocentric frame of $\sim 1,100$ RR Lyrae stars located within a Galactocentric radius of 50-100 kpc with those from the mock star catalogs generated by our 50 MW-LMC simulations. We observe that the RR Lyrae sample exhibits a mean latitudinal velocity of $\langle v_{b}\rangle=18.1\pm 4.1$ km/s.
  • This result aligns with the outcomes ($\langle v_{b}\rangle\approx 16\pm 4$ km/s) from simulations in which the LMC follows a first-infall orbital scenario and the MW halo is spherical ($q=1.0$). In contrast, the results from simulations where the LMC has a second-passage orbital scenario are consistently lower than $8\pm 4$ km/s for various parameter combinations. Therefore, our comparative study supports the first-infall scenario for the LMC.